# https://replicantlife.com

### 09 Jan 2024 v1

@@
Jason Toy Jason Toy, somatic.io - jasontoy@gmail.com
Phil Tabor
Josh MacAdam, obvi
     Josh MacAdam    and Phil Tabor
@@ 3 Architecture
Our memory recall system can be seen as an agent optimized version of Retrieval-Augmented Generation (RAG) [8]. Anytime an agent makes a decision, relevant memories are retrieved to prime the agent on what to do. There are several memory types such as observance memories like “John saw a cat” and conversation memories like “John said ’How are you doing Paul?”. We explicitly store metacognition memories where an agent looks at its past memories and actions and asks a meta-question. That thought is stored as a meta-memory and inserted into the memory stream of the agent. In future actions and conversations, these meta-memories are recalled along with other memories to prime the agent to think about these meta-thoughts when in conversation and taking action. For each step in the simulation, we allow the agent to choose an action from a list of possible actions which also includes the meta_cognize function.
@@ 4 Simulation Framework
To conduct these experiments we built a framework dubbed ReplicantLife, where agents can be run standalone or within a simulated town environment. ReplicantLife has a pluggable architecture which allows can utilize any popular LLM with an http interface, including locally hosted models through ollama. There is preliminary support for concurrency using threads. Concurrency is limited by GPUs and LLM calls, so max concurrency should be set to the total available GPUs. Feature flags are provided to toggle various functionality including LLM call limits.
@@ 5 Experiments
We tested our simulation framework in a variety of different situations including a Christmas party, zombie apocalypse, and murder mystery. In the Christmas party simulation agents hosted a party where multiple other agents were invited and arrived at the specified time to attend. This is similar to the earlier work in generative agents[3] where agents had to coordinate a similar social activity.
@@
To measure the performance, we opted not to use full human evaluations due to resources. Instead we opted to use LLMS to assist us in evaluating performance, a technique similary described as LLM-as-a-Judge in [11]. In their paper, they found that using an LLM to judge evaluations is 80% in agreement with human judges. We found in our own spot checking of evaluations that LLM performance was just as good as a human judge. The majority of our tests were run using Mistral 7B[10], but we also did extensive testing with Phi1[12], Phi2[13] Llama2[14], Mixtral [15], GPT-3.5-turbo, GPT4 [16], and other models. We standardized on Mistral due to its combination of speed, small model size, and excellent performance. While we did build support for ChatGPT models, we primarily focused on local LLMs for cost and performance reasons. We built out test infrastucture to spin up LLM nodes on public GPU clouds when needed for faster simulations.
@@ 6.2 Future Directions
With our current metacognition implementation, we built a base framework that shows increased performance of generative agents to achieve their tasks. We believe our architect can be improved along several dimensions.
@@ 6.2.1 Improved memory retrieval
In high dimensional spaces, vectors that are similar may not be sementically related. Returning irrelavent memories would effect the output of the LLM and in turn the actions the agent takes. Cosine similarity is what is used in almost all RAG[8] implementations and so many of these systems will potentially have similar issues of non-relevant context being included. Changing out embedding models could improve performance, but the cosine similarity issue would still remain. As recently adopted in [18], improvements can be made to relevance scoring by searching a database based on the LLM’s output embedding and employing a K-nearest neighbors (KNN) search algorithm. This process selectively adjusts the output embedding vector prior to token generation. This reverse sequence of operations, wherein LLM embeddings are utilized, has demonstrated superior performance compared to existing methods like RAG, as reported in the literature. Retrieving more relevant memories would likely improve performance in all elements of the system. Sometimes when someone has a thought about a topic, other seemingly unrelated thoughts may appear. So in some sense, this may not be a big problem. Further experiments would have us test different memory augmentation and retrieval models.
@@ 6.2.2 Inference Speed
We also explored different inference engines. We tested PowerInfer [20], an inference engine that exploits high locality in LLM inference. The authors convert common models that use ReLU into another format that uses another predictive model to read the input query and selectively choose which neurons to activate. In effect, this reduces the total amount of neurons and computation needed to process a prompt. We did not see speed improvements on the models we tested with. We also tested with vLLM from [21], an inference engine that uses PagedAttention and saw a 35% speed increase when using concurrency, but further testing is needed. We would want to continue testing with other models, different concurrency systems, and inference speed up techniques to get large simulations to run in realtime. Large groups of simulated agents on non-cloud based machines could be interesting for sandbox games and simulations.
@@ 6.2.5 Metacognition directly in a LLM
Our current implementation of metacognition uses Python to essentially graft on metacogntion on top of an LLM. Future investigations may delve into building metacognition like capabilities directly into the LLM, allowing them to introspect and enhance their own decision-making processes. This introspective capability could contribute potentially paving the way for more adaptive and self-aware systems. One interesting view of the human brain is as a coordinate transformation engine: “A brain is a well-designed machine for the frame conversion to internalize the external world” [23], “the egocentric representations of the primary sensory cortical areas must be transformed into an allocentric representation in the hippocampus, and then transformed back to an egocentric motor representation for behavioral output” [24], and “Our findings provide compelling evidence that the reference frame of neural representations is not static and can be powerfully modulated by task instructions.” [25]. The human brain has been found to store over 10 different coordinate representations along with orientations or points of view, such as allocentric and egocentric orientations.
@@
Our hypothesis posits that by adapting the underlying architecture to accommodate more dynamic representation transformations, a neural network can be trained to facilitate metacognitive processes. Achieving this would likely involve formulating a hybrid objective function wherein the model learns to not only predict the next token, but also to evaluate the token’s quality in relation to the input query.

### 29 Feb 2024 v2

@@
Jason Toy Jason Toy, replicantlife.com - jasontoy@gmail.com    Phil Tabor Phil Tabor, neuralnet.ai - phil@neuralnet.ai    Josh MacAdam
@@
Figure 2: A timeline showing an agent’s goals and thoughts change as it interacts in the simulation.
@@ 3 Architecture
Retrieval-Augmented Generation (RAG) [8] is a technique created to give LLMs the ability to surface and reference content that the model was not explicitly trained on. Our memory recall system can be seen as a dynamic RAG system. As the agent’s system acquires new memories and metacognizes over time, the relevant memories sent to the LLM change.
Anytime an agent makes a decision, relevant memories are retrieved to prime the agent on what to do. There are several memory types such as observance memories like “John saw a cat” and conversation memories like “John said ’How are you doing Paul?”. We explicitly store metacognition memories where an agent looks at its past memories and actions and asks a meta-question. That thought is stored as a meta-memory and inserted into the memory stream of the agent. In future actions and conversations, these meta-memories are recalled along with other memories to prime the agent to think about these meta-thoughts when in conversation and taking action. For each step in the simulation, we allow the agent to choose an action from a list of possible actions which also includes the meta_cognize function.
@@ 4 Simulation Framework
To conduct these experiments we built a framework dubbed ReplicantLife, where agents can be run standalone or within a simulated town environment. ReplicantLife has a pluggable architecture which can utilize any popular LLM with an http interface, including locally hosted models through ollama. There is preliminary support for concurrency using threads. Concurrency is limited by GPUs and LLM calls, so max concurrency should be set to the total available GPUs. Feature flags are provided to toggle various functionality including LLM call limits.
@@ 5 Experiments
We tested our simulation framework in a variety of different situations including a Christmas party, zombie apocalypse, and murder mystery. In the Christmas party simulation agents hosted a party where multiple other agents were invited and arrived at the specified time to attend. This is similar to the earlier work in generative agents[3] where agents had to coordinate a social activity.
@@
To measure performance, we opted not to use full human evaluations due to resource constraints. Instead we opted to use LLMS to assist us in evaluating performance, a technique similary described as LLM-as-a-Judge in [11]. In their paper, they found that using an LLM to judge evaluations is 80% in agreement with human judges. We found in our own spot checking of evaluations that LLM performance was just as good as a human judge. The majority of our tests were run using Mistral 7B[10], but we also did extensive testing with Phi1[12], Phi2[13] Llama2[14], Mixtral [15], GPT-3.5-turbo, GPT4 [16], and other models. We standardized on Mistral due to its combination of speed, small model size, and excellent performance. While we did build support for ChatGPT models, we primarily focused on local LLMs for cost and performance reasons. We built out test infrastucture to spin up LLM nodes on public GPU clouds when needed for faster simulations.
@@ 6.2 Future Directions
With our current metacognition implementation, we built a base framework that shows increased performance of generative agents to achieve their tasks. We believe our architecture can be improved along several dimensions.
@@ 6.2.1 Improved memory retrieval
In high dimensional spaces, vectors that are similar may not be sementically related. Returning irrelavent memories would effect the output of the LLM and in turn the actions the agent takes. Cosine similarity is what is used in most RAG[8] implementations and so many of these systems will potentially have similar issues of non-relevant context being included. Changing out embedding models could improve performance, but the cosine similarity issue would still remain. As recently adopted in [18], improvements can be made to relevance scoring by searching a database based on the LLM’s output embedding and employing a K-nearest neighbors (KNN) search algorithm. This process selectively adjusts the output embedding vector prior to token generation. This reverse sequence of operations, wherein LLM embeddings are utilized, has demonstrated superior performance compared to existing methods like RAG, as reported in the literature. Retrieving more relevant memories would likely improve performance in all elements of the system. Further experiments would have us test different memory augmentation and retrieval models.
@@ 6.2.2 Inference Speed
We also explored different inference engines. We tested PowerInfer from textcitesong2023powerinfer, an inference engine that exploits high locality in LLM inference. The authors convert common models that use ReLU into format that uses another predictive model to read the input query and selectively choose which neurons to activate. In effect, this reduces the total amount of neurons and computation needed to process a prompt. We did not see speed improvements on the models we tested with. We also tested with vLLM from [20], an inference engine that uses PagedAttention and saw a 35% speed increase when using concurrency, but further testing is needed. We would want to continue testing with other models, different concurrency systems, and inference speed up techniques to get large simulations to run in realtime. Large groups of simulated agents on non-cloud based machines could be interesting for sandbox games and simulations.
@@ 6.2.5 Metacognition directly in a LLM
Our current implementation of metacognition uses Python to essentially graft on metacogntion on top of an LLM. Future investigations may delve into building metacognition like capabilities directly into the LLM, allowing them to introspect and enhance their own decision-making processes. This introspective capability could contribute potentially pave the way for more adaptive and self-aware systems. One interesting view of the human brain is as a coordinate transformation engine: “A brain is a well-designed machine for the frame conversion to internalize the external world” [22], “the egocentric representations of the primary sensory cortical areas must be transformed into an allocentric representation in the hippocampus, and then transformed back to an egocentric motor representation for behavioral output” [23], “Our findings provide compelling evidence that the reference frame of neural representations is not static and can be powerfully modulated by task instructions.” [24], “Conjunctive representations among input variables appear in many theoretical models for neural systems that perform coordinate transformations” [25], and “This picture also implies operation of different representations over different timescales… a process of translation between the systems” [26]. The human brain has been found to store over 10 different coordinate representations along with orientations or points of view, such as allocentric and egocentric orientations.
A large portion of computational neuroscience studies are devoted to decoding how the brain transforms and integrates between the myriad of neural representations.
@@
Our hypothesis posits that by adapting the underlying architecture to accommodate more dynamic representation transformations, a neural network can be trained to represent a wider range of data and facilitate metacognitive processes. Achieving this would likely involve formulating a hybrid objective function wherein the model learns to not only predict the next token, but also to evaluate the token’s quality in relation to the input query.